Circular No. 8294

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 D1 (NEAT) An apparently asteroidal object reported by the NEAT project (dis-

covery observation below), and posted on the NEO Confirmation Page, has been found to show apparent cometary appearance on CCD images taken by J. Tichá and M. Tichý (Kleť, Feb. 20.8 UT; diffuse with a faint coma of diameter 10′′ and a possible tail in p.a. 225^{o}) and by P. Birtwhistle (Great Shefford, Berkshire, U.K., Feb. 22.8 and 23.8; somewhat larger and less concentrated than stars of similar brightness, with a coma diameter of 9′′–10′′, slightly extended in p.a. 270^{o}).

2004 UT  α_{2000}  δ_{2000}  Mag.
Feb. 17.12122  3^{h}42^{m}57.^{s}50  +24^{o}35′15′′.4  19.8

Additional astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2004-D44.

T = 2006 Feb. 11.605 TT  ω = 75.^{o}064
Ω = 62.234 }2000.0
q = 5.01497 AU  i = 45.544

SUPERNOVA 2004Z IN MCG +10-19-85 R. Ellis, California Institute of Technology (CIT); M. Sullivan, Univer-

sity of Toronto; P. Nugent, Lawrence Berkeley National Laboratory; and A. Gal-Yam, CIT, report that a CCD spectrum (range 360–950 nm) of SN 2004Z (cf. IAUC 8290), obtained on Feb. 22 UT with the Keck-I 10-m reflector (+ LRIS), reveals it to be a relatively young type-II supernova, featuring a blue continuum with strong P-Cyg profiles of the hydrogen Balmer lines as well as He I (587.6 nm). The maximum velocity of Hα in SN 2004Z is ≈ 10000 km/s.

BZ URSAE MAJORIS E. Waagen, AAVSO, reports an outburst of BZ UMa, the first since

Dec. 2002 (when it reached m_{v} = 11.4 on Dec. 19 before declining to mag 14.5 by Dec. 23), as indicated by the following magnitude estimates, visual unless otherwise noted: Feb. 23.922 UT, [14.0 (E. Muyllaert, Oostende, Belgium); 24.138, V = 16.2 (W. MacDonald, Winchester, Ontario, Canada; CCD); 25.294, 11.6 (M. Simonsen, Imlay City, MI); 25.813, 11.0 (H. McGee, Guildford, England); 26.209, 11.5 (R. Smith, W. Kingsdown, England).

2004 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green